find the slope of the line that crosses the x-axis at -6 and the y-axis -2.
Answer by checkley71(8403)

THESE POINTS ARE: (-6,0) & (0,-2)
THUS THE SLOPE IS: (-2-0)/(0+6)=-2/6=-1/3 FOR THE SLOPE: